two points A and B lie on a radial line of a rotating disk.The points are 50 mm apart.Va=214m/min and Vb=270m/min.Determine the radius of rotation for each of these points


1
Expert's answer
2022-01-20T08:11:02-0500

Let "r_A=" the radius of rotation for point "A" and let "r_B=" the radius of rotation for point "B."

Given "r_B-r_A=50\ mm=0.05\ m."

The angular velocity "\omega" will be same. Then


"v_A=\omega r_A, v_B=\omega r_B"

"\dfrac{v_A}{r_A}=\dfrac{v_B}{r_B}"

"r_B=\dfrac{v_B}{v_A}r_A"

Then


"r_A+0.05=\dfrac{270\ m/min}{214\ m/min}\cdot r_A"

"(270-214)r_A=10.7"

"r_A=\dfrac{10.7}{56}\ m"

"r_A=0.191\ m=191\ mm"

"r_B=0.191\ m+0.05\ m=0.241\ m=241\ mm"


"r_A=191\ mm"

"r_B=241\ mm"